OpenStudy (anonymous):

is -1/9 greater then -2/9

OpenStudy (anonymous):

i think so..........

OpenStudy (studygurl14):

Yes. With negative numbers, the seemingly "smaller" number is actually bigger. \(\large +\frac{1}{9}<+\frac{2}{9}\) \(\large -\frac{1}{9}>-\frac{2}{9}\)

OpenStudy (michele_laino):

yes, because 1/9 is less than 2/9

OpenStudy (michele_laino):